# Well behaved class of Heintzmann’s solution within $f(R,\,T)$ framework

### v1

@@ V.2 Nature of pressure and density
Also, we can calculate the central pressure and central density as,
@@
Refer to captionRefer to caption
Figure 3: The pressure and density gradients are shown against ‘r’.
Refer to captionRefer to caption
Figure 4: $\frac{d^{2}\rho}{dr^{2}}$ and $\frac{d^{2}p}{dr^{2}}$ are shown against ‘r’.
@@ V.3 Energy Conditions
Since the solutions to the EFEs indicate the presence of ordinary and exotic matter within material objects, it is anticipated that any solution, whether in improved gravity or Einstein gravity, must meet all energy criteria to show the existence of material objects. The null energy condition (NEC), the weak energy condition (WEC), the strong energy condition (SEC), and the dominant energy condition (DEC) are characterized as energy restrictions and are crucial to investigate the existence of a realistic matter distribution. Depending on the energy-momentum tensor $T_{\mu\nu}$ the Energy conditions are described as:
@@
Here, $k^{\nu}$ is a null vector and $U^{\mu}$ is a time-like vector. In this connection, it is to be mentioned that where $U^{\mu}$ is a time-like vector but $T_{\mu\nu}U^{\mu}$ not spacelike.
So, we will need the following expressions to validate the energy criteria given earlier.
@@ V.5 Relativistic adiabatic index
The adiabatic index, which is used to describe the stiffness of the EoS for a given energy density, is crucial when talking about the stability of both relativistic and non-relativistic compact objects Sharifa and Waseemb (2018). According to the assessments of some researchers, within a dynamically stable stellar object, the adiabatic index value is expected to be greater than $\frac{4}{3}$ Chandrasekhar (1964); Heintzmann and Hillebrandt (1975); Hillebrandt and Steinmetz (1976). The relativistic adiabatic index $\Gamma$ can be expressed as follows:
$\Gamma$ $=$ $\frac{\rho+p}{p}V^{2},$ (39)
@@ V.6 Modified TOV Equation
Any celestial object experiences equilibrium, or dynamical balance, when many internal forces act simultaneously. This condition is characterized by the generalized Tolman-Oppenheimer-Volkoff (TOV) equation. For our current isotropic stellar model in the $f(R,T)$ gravity system, the equilibrium condition is given by the following modified TOV equation.
@@
Ultimately, FIG. 7 allows us to conclude that our present model remains in stable equilibrium because the gravitational force counterbalances the combined behavior of hydrostatic force and modified gravity force.
@@ V.7 Equation of state (EoS)
An essential astrophysical tool that could help us comprehend the basic principles of matter dispersion is the EoS parameter. The EoS illustrates how matter is influenced by a suggested set of physical specifications. The relationship between pressure and matter density is simply explained by the equation of state. While several researchers used linear, quadratic, polytropic, and other EoS to represent the compact object. But we did not make any particular assumptions about EoS in this work. Consequently, $\omega=p/\rho$ is denoted as EoS parameter, which is commonly represented as a dimensionless number that may be used to characterize the relationship between matter density and pressure.
Refer to captionRefer to caption
@@
Its different values correspond to different stages of the entire cosmos. In terms of Zeldovich’s criterion, any fluid sphere with a positive pressure-to-density ratio smaller than unity is regarded as physically acceptable Zel’dovich (1962); Zeldovich and Novikov (1971). Accordingly, $\omega$ must always be continuous and fall between 0 and 1 throughout the stellar medium and this is satisfied by our solutions. Moreover, it provides the radiation phase when $\omega$ is between 0 and 1 and so describes also the radiating nature of our proposed compact objects. We have used a parametric plot to show the changes in pressure with respect to matter density in FIG. 8. For various values of $\chi$, the variation in pressure-to-density ratio ($\omega$) is also shown in FIG. 8.
@@
V.8 Effective mass-radius relationship, effective compactness and effective surface redshift
The effective mass function of the present system can be found as :
@@ V.8 Effective mass-radius relationship, effective compactness and effective surface redshift
It is simple to verify that the effective mass of a compact star is simply proportional to its radius, but it depends on the parameter $\chi$. The effective mass function at the core is regular, as seen in FIG. 9. Also, we can see that the maximum mass is attained at the boundary of the star. The effective compactness factor ($u^{\text{eff}}$) is expressed as:
$u^{\text{eff}}(r)=\frac{m^{\text{eff}}(r)}{r}$ (46)
assigns the compact objects into the following groups: normal star ($u^{\text{eff}}\sim 10^{-5}$), white dwarfs ($u^{\text{eff}}~{}\sim 10^{-3}$), neutron star ($410^{-1}<~{}u^{\text{eff}}<1/$ 4), ultra-compact star ($21/4<~{}u^{\text{eff}}<1/$ 2), and black hole ($u^{\text{eff}}\sim 1/2$). Moreover, the effective surface redshift ($z_{s}^{\text{eff}}$) can be computed using the formula below:
$z_{s}^{\text{eff}}(r)$ $=$ $\frac{1}{\sqrt{1-2u^{\text{eff}}}}-1,$ (47)
Table 4: Numerically estimated values of effective mass, effective compactness factor, and effective surface redshift at the surface $r=\mathfrak{R}$ correspond to the compact star Vela X-1 for different values of coupling constant $\chi$ (Taking $B=0.001~{}\rm{km}^{-2}$).
$\chi$ $m^{\text{eff}}(\mathfrak{R})$ $u^{\text{eff}}(\mathfrak{R})$ $z_{s}^{\text{eff}}(\mathfrak{R})$
@@
Refer to captionRefer to captionRefer to caption
Figure 9: (left) The effective mass, (middle) the effective compactness and (right) the effective surface redshift are plotted against the radius inside the stellar interior.
FIG. 9 displays the profiles of effective compactness and effective surface redshift. It is shown that they are all monotonically increasing with respect to $r$ confirming the viability of our presented model.
@@ VI Discussion
  • FIG. 9 displays the variations of the effective mass function, effective compactness factor, and effective surface redshift for different values of $\chi$ within compact object formations. From figure, we see that they are all well-behaved, increasing steadily with $r$ and attain their peak values at the boundary. We have estimated the values of effective mass, effective compactness factor, and effective surface redshift at the stellar surface in Table 4.

### v2

@@ V.2 Nature of pressure and density
Also, we can calculate the central pressure and central density under modified gravity as
@@
Now we are interested in evaluating the proper range of the coupling parameter $\chi$.
Case I
We initially assume that, $\chi+2\pi>0$.
Now $p_{c}>0\implies\chi>-\frac{2(3-C)\pi}{3}$.
So we obtain,
$\chi>Max\Bigg{\{}-2\pi,-\frac{2(3-C)\pi}{3}\Bigg{\}}.$ (35)
Also since, $\rho_{c}-p_{c}>0\implies\frac{3BC}{2\chi+4\pi}>0\implies C>0.$
Again, $\rho_{c}>0\implies\chi>-\frac{6(1+C)\pi}{2+3C}$. So, it is clear that $\Bigg{[}-\frac{2(3-C)\pi}{3}-\Bigg{\{}-\frac{6(1+C)\pi}{2+3C}\Bigg{\}}\Bigg{]}>0$, for $C>0$. So, with the help of (35) and $C>0$, again we can repeat, $\chi>Max\Bigg{\{}-2\pi,-\frac{2(3-C)\pi}{3}\Bigg{\}}$.
Also, $C>0\implies-\frac{2(3-C)\pi}{3}>-2\pi$.
Hence, we can finally obtain,
$\chi>-\frac{2(3-C)\pi}{3}.$ (36)
Case II
Let us assume, $\chi+4\pi<0\implies\chi+2\pi<0$ also.
Now, $\rho_{c}-p_{c}>0\implies C<0$. Again, $p_{c}>0\implies\chi>-\frac{2(3-C)\pi}{3}$.
So, with the help of $C<0$ we obtain,
$-\frac{2(3-C)\pi}{3}<\chi<-2\pi.$ (37)
Case III
Now let us assume, $-4\pi<\chi<-2\pi$.
Now, $p_{c}>0\implies\chi<-\frac{2(3-C)\pi}{3}$.
Now, $\rho_{c}-p_{c}>0\implies C<0.$ So, obviously $-\frac{2(3-C)\pi}{3}<-2\pi$.
Hence, we obtain the range for $\chi$ as,
$-4\pi<\chi<Min\Bigg{\{}-2\pi,-\frac{2(3-C)\pi}{3}\Bigg{\}}.$ (38)
@@
Figure 3: The density and pressure gradients ($\frac{d\rho}{dr}$ and $\frac{dp}{dr}$) are plotted against ‘r’.
Figure 4: Second order derivatives of density and pressure ($\frac{d^{2}\rho}{dr^{2}}$ and $\frac{d^{2}p}{dr^{2}}$) are plotted against ‘r’.
@@ V.3 Energy Conditions
Pointwise energy conditions were first proposed as physically plausible constraints on matter within the framework of mathematical relativity. For example, they seek to convey the attractiveness of gravity or the positivity of mass. More significantly, they have served as presumptions in mathematical relativity to support the non-existence of wormholes and other exotic phenomena as well as singularity theorems. Energy conditions play a significant impact in the context of the exotic properties. These conditions restrict the contraction of the stress tensor at each spacetime point. The null energy condition (NEC), the weak energy condition (WEC), the strong energy condition (SEC), and the dominant energy condition (DEC) are characterized as the four main energy restrictions and are crucial to investigate the existence of a realistic matter distribution. To rewrite these energy conditions in Einstein’s equation, we can replace the stress tensor with the Ricci curvature tensor, resulting in a geometric form rather than the original physical form. Depending on the energy-momentum tensor $T_{\mu\nu}$, the physical forms of these Energy conditions are described as:
@@
Here, $k^{\nu}$ is a null vector and $U^{\mu}$ is a time-like vector. In this connection, it is to be mentioned that where $U^{\mu}$ is a time-like vector but $T_{\mu\nu}U^{\mu}$ is not space-like.
So, we will need the following expressions to validate the energy conditions given earlier.
@@ V.5 Relativistic adiabatic index
The adiabatic index, which is used to describe the stiffness of the EoS for a given energy density, is crucial when talking about the stability of both relativistic and non-relativistic compact objects Sharifa and Waseemb (2018). In order to derive the equations governing the radial oscillations both spacetime and fluid variables are perturbed so as not to violate the spherical symmetry of the background body. In the unperturbed model, if we define $\delta r(r,t)$ as the time-dependent radial displacement of a fluid element positioned at location $r$ and assuming a harmonic time dependence-
$\delta r(r,t)=X(r)e^{i\omega t}$
($\omega$ be the frequency of the oscillation), we derive the subsequent formula that characterizes the radial oscillations as in Ref. Kokkotas and Ruoff (2001):
$V^{2}X^{\prime\prime}+[(V^{2})^{\prime}-Z+4\pi r\gamma pe^{2\lambda}-\nu^{\prime}]X^{\prime}+[2(\nu^{\prime})^{2}+\frac{2m}{r^{3}}-Z^{\prime}-4\pi(\rho+p)Zre^{2\lambda}+\omega^{2}e^{2\lambda-2\nu}]X=0.$
Where $V$ is the sound speed, which is calculated from the unperturbed background for a specific EoS,
$V^{2}=\frac{dp}{d\rho}$
and $\Gamma$ is the adiabatic index, which is connected to the sound speed through
$\Gamma$ $=$ $\frac{\rho+p}{p}\frac{dp}{d\rho},$ (45)
@@
According to the assessments of some researchers, within a dynamically stable stellar object, the adiabatic index value is expected to be greater than $\frac{4}{3}$ Chandrasekhar (1964); Heintzmann and Hillebrandt (1975); Hillebrandt and Steinmetz (1976).
@@ V.6 Modified TOV Equation
Every celestial object remains in static equilibrium, or dynamical balance, under different internal forces acting simultaneously on it. This condition is characterized by the generalized Tolman-Oppenheimer-Volkoff (TOV) equation. For our current isotropic stellar model in the $f(R,T)$ gravity system, the equilibrium condition is described by the following modified TOV equation.
@@
Ultimately, FIG. 7 evidently allows us to conclude that our present model attains its static equilibrium under these three different forces because the gravitational force counterbalances (or nullifies) the combined effect of hydrostatic force and modified gravity force.
Although, the static equilibrium in consideration may exhibit stability or instability in response to a minor radial perturbation. Therefore, in the last subsection (V.9), we have conducted the Harrison-Zeldovich-Novikov stability test in order to further investigate whether this static equilibrium is stable or unstable.
@@ V.7 Equation of state (EoS)
An essential astrophysical tool that could help us comprehend the basic principles of matter dispersion is the EoS parameter. The EoS illustrates how matter is influenced by a suggested set of physical specifications. The relationship between pressure and matter density is simply explained by the equation of state. While several researchers used linear, quadratic, polytropic, and other EoS to represent the compact object. However, we did not make any particular assumptions about EoS in this work. Consequently, $\omega=p/\rho$ is denoted as the EoS parameter, which is commonly represented as a dimensionless number that may be used to characterize the relationship between matter density and pressure.
@@
Its different values correspond to different stages of the entire cosmos. Regarding the stability of compact stellar objects, the Zeldovich requirement for stability is generally valid for a wide range of gravity theories, including modified ones. Several investigations use the Zeldovich condition for stability in modified gravity theories Bhar et al. (2019). Even though it is helpful, modified gravity theories cannot be guaranteed to be stable by the Zeldovich criterion alone. It gives a fundamental criterion based on gravitational and thermal energies but ignores additional aspects such as magnetic fields, rotation, and the detailed dynamics of gravity in modified theories. The gravitational potential and equations of motion in modified gravity are not the same as those in general relativity. The modified gravity systems introduce additional terms that can significantly alter the stability conditions and can lead to deviations from the standard Zeldovich stability criterion Jasim et al. (2021). In order to completely comprehend stability in these scenarios, it is necessary to go further into the particular theory, even though the Zeldovich condition offers a starting point. In terms of Zeldovich’s criterion, any fluid sphere with a positive pressure-to-density ratio smaller than unity is regarded as physically acceptable Zel’dovich (1962); Zeldovich and Novikov (1971). Accordingly, $\omega$ must always be continuous and fall between 0 and 1 throughout the stellar medium and this is satisfied by our solutions. Moreover, $\omega=1/3$ represents an equation of state (EoS) for radiation for the matter constituent being relativistic Das et al. (2016). But in this study, we get several EoS parameter $\omega$ values other than $\omega=1/3$. This suggests that our model is more appropriate for compact stars that radiate. We have used a parametric plot to show the changes in pressure with respect to matter density in FIG. 8. For various values of $\chi$, the variation in pressure-to-density ratio ($\omega$) is also shown in FIG. 8.
@@
V.8 Mass-radius relationship, compactness and surface redshift
The mass function of the present system can be found as :
@@ V.8 Mass-radius relationship, compactness and surface redshift
The mass function at the core is regular, as seen in FIG. 9. It is simple to verify from FIG. 9 that the mass function of a compact star is simply proportional to its radius, but it depends on the parameter $\chi$. Also, we can see that the maximum mass is attained at the boundary of the star. The compactness factor ($u$) is expressed as:
$u(r)=\frac{m(r)}{r}$ (52)
assigns the compact objects into the following groups: normal star ($u\sim 10^{-5}$), white dwarfs ($u~{}\sim 10^{-3}$), neutron star ($410^{-1}<~{}u<1/$ 4), ultra-compact star ($21/4<~{}u<1/$ 2), and black hole ($u\sim 1/2$). Moreover, the surface redshift ($z_{s}$) can be computed using the formula below:
$z_{s}(r)$ $=$ $\frac{1}{\sqrt{1-2u}}-1,$ (53)
Table 4: Numerically estimated values of mass, compactness factor, and surface redshift at the surface $r=\mathfrak{R}$ correspond to the compact star Vela X-1 for different values of coupling constant $\chi$ (Taking $B=0.001~{}\rm{km}^{-2}$).
$\chi$ $m(\mathfrak{R})$ $u(\mathfrak{R})$ $z_{s}(\mathfrak{R})$
@@
Figure 9: (left) The mass function, (middle) the compactness and (right) the surface redshift are plotted against the radius inside the stellar interior.
FIG. 9 displays the profiles of compactness and surface redshift. It is shown that they are all monotonically increasing with respect to $r$ confirming the viability of our presented model.
V.9 Stability Test through Harrison-Zeldovich-Novikov criterion
We are particularly interested in studying the stability of the compact stellar configuration. The Harrison-Zeldovich-Novikov criterion Harrison et al. (1965); Zeldovich and Novikov (1971) describes the stability of a compact object. The necessary (but not sufficient) condition for the stability of a compact stellar configuration is that the total gravitational mass ($M$) of a compact star should be an increasing function of the central density ($\rho_{c}$) throughout the stellar region. Mathematically defined as, $\frac{\partial M}{\partial\rho_{c}}>0$ Glendenning and Kettner (2000); Arbañil and Malheiro (2016). This condition, well beyond the purview of GR, has been commonly used in the literature of modified gravity theories also because of its simplicity. The dependency of compact star’s total mass in gravitational units on their central density is depicted in FIG. 10 and we can verify that the total mass ($M(\rho_{c})$) increases with increasing central density ($\rho_{c}$). Also the right panel of FIG. 10 describes the positive behaviour of $\frac{\partial M}{\partial\rho_{c}}$ with respect to central density ($\rho_{c}$). Both curves exhibit asymptotic behaviour. Thus we can conclude that our current model satisfies Harrison-Zeldovich-Novikov stability criterion and maintains stable equilibrium throughout the region.
Figure 10: (left) Total mass versus central density and (right) $\frac{\partial M}{\partial\rho_{c}}$ versus central density plots for the chosen compact stellar configuration.
@@ VI Discussion
  • FIG. 9 displays the variations of the mass function, compactness factor, and surface redshift for different values of $\chi$ within compact object formations. From figure, we see that they are all well-behaved, increasing steadily with $r$ and attain their peak values at the boundary. We have estimated the values of mass function, compactness factor, and surface redshift at the stellar surface in Table 4.
  • The left panel of FIG. 10 shows how the total mass varies with central density. While the right panel exhibits the feature $\frac{\partial M}{\partial\rho_{c}}>0$ for our system and demonstrates system stability by adhering to the Harrison-Zeldovich-Novikov stability criterion.